Given the following revenue and cost functions:

Revenue = -4x2+199x

Cost = 44x+7

Find the highest breakeven point
Answer #1

At the breakeven point revenue = total cost

Thus, - 4x2 + 199x = 44x + 7

Which gets us to the equation

-4x2 + 155x - 7 =0

Solving Steps - 4x² + 155x - 7 = 0 Change the signs 4x2 - 155x + 7 = 0 Solve the quadratic equation -(-155) + 7 (-155) 2-4x4x

X depicts the number of units

Since the value of x1 is very small we assume the quantity of units be 39 (rounded off)

Thus 39 is the highest break even pont.